# excitations, and scale dependence

### v1

@@ 1 Introduction
Under the inflationary paradigm, the Universe underwent a rapid phase of expansion in its earlier moments (Guth 1981; Linde 1982; Guth 2004). This process also generates the initial fluctuations in the density field that subsequently grow and are observed in the cosmic microwave background and the large-scale structure. The fluctuations are commonly modeled to follow the statistics of a Gaussian random field. However, deviations from Gaussianity—called “Primordial Non-Gaussianities” (PNGs)—can arise from particle interactions during inflation (see Chen 2010a, for a review). Detecting such PNGs would provide a unique window into the microphysics of the early Universe, offering direct clues about the field content and interactions operative during inflation.
@@
There are also extensions to these standard scenarios that capture interactions of the inflaton with additional massless/massive particles. This framework, known as “cosmological collider physics” (Chen & Wang 2010; Baumann & Green 2012; Arkani-Hamed & Maldacena 2015; Lee et al. 2016; Arkani-Hamed et al. 2020), has recently been studied in the large-scale structure context by Goldstein et al. (2024) for the squeezed-limit bispectrum and by Anbajagane & Lee (2025, henceforth, Paper I) for the full bispectrum. In particular, in Paper I we introduced a method for generating initial density fields with arbitrary primordial bispectra. This enabled us to resolve, for the first time, over thirty collider PNG templates in the nonlinear regime. The inflationary signatures in structure formation from this regime are distinct from those signatures found on linear and quasi-linear scales, and therefore provide a complementary avenue for placing constraints on inflationary models (Anbajagane et al. 2023, Paper I). Furthermore, Paper I showed that the potential lensing-only constraints on such collider models can be competitive and complementary to the existing constraints from Planck 2018 (Sohn et al. 2024).
@@ 2.1 Generating Initial Conditions from Arbitrary Bispectra
To generate initial conditions with a given bispectrum, we follow the approach of Scoccimarro et al. (2012). This method requires the bispectrum (or kernel) to be separable, i.e., $B(k_{1},k_{2},k_{3})=f(k_{1})g(k_{2})h(k_{3})$. This condition is not always met for all bispectra considered in this work. Instead, we follow the approach of Paper I by decomposing the bispectra into a sum of terms that are each individually factorizable as
@@ 2.2 Inflationary Models
We consider a wide variety of models following those explored in Planck Collaboration et al. (2014, 2016, 2020b), and choose a subset to simulate in our work. Many of these models, such as the scale-dependent ones or linear resonances, are manifestly separable (Münchmeyer et al. 2014). Others that are not manifestly separable due to a factor of $1/(k_{1}+k_{2}+k_{3})^{n}$, with $n>0$, can be approximated using the Schwinger parameterization (e.g., Smith & Zaldarriaga 2011). Thus, many bispectra also have template-specific approaches to decomposing them. Instead of following different methods tailored to each given bispectrum, we follow the Planck analyses in using mode decomposition methods on all templates, given we scan a wide range of model space. We now detail our chosen models below.
@@
We then consider templates corresponding to resonant particle interactions. In this case, the inflaton potential has a localized feature that generates oscillations in the bispectrum (see Achúcarro et al. 2022, for a review). The models used in Paper I also exhibit oscillations, but are derived from explicit couplings of the inflaton to other massive fields (e.g., Arkani-Hamed & Maldacena 2015), whereas the models we study here are described by phenomenological parameters alone. Sharp changes to the inflaton potential also induce oscillations at other orders, such as the power spectrum and trispectrum. In this work, we will first focus on oscillations only in the bispectrum, but then consider a class of models where the bispectrum and power spectrum both contain oscillations. We ignore the primordial trispectra as its impact on late-time structure is subdominant to the power spectra and bispectra. [^1] Note that there do exist some models where the primordial trispectrum can dominate over the bispectrum (Philcox 2025a, b). In general, the derivative of a signature in late-time structure, $X$, with the bispectrum amplitude $f_{\rm NL}$ is $\mathcal{O}(10^{3})$ higher than that with the trispectrum amplitude, $g_{\rm NL}$ (e.g., LoVerde & Smith 2011, see their ).
Oscillatory bispectra can be broadly classified into linear and logarithmic templates, given by (Chen et al. 2007b; Chen 2010a) {mybox}
@@
where $\omega$ is the frequency of the oscillations and $\phi=0$ is the phase. We explore a fairly small range in $\omega\in[0.1,2.5]$ relative to perturbation theory-based analyses of the CMB (e.g., Beutler et al. 2019a). Our choice is limited by the basis functions used in decomposing the templates. Oscillations of high order (large $w$) are not accurately reproduced by our existing basis functions. [^2] In principle, this can be alleviated by using an adequately high order for the Legendre polynomials in Eq. (2.3). However, doing so degrades the orthogonality of the basis set as discussed in Section 2.1 of Paper I. Higher values of $w$ could be explored by making improvements to the choice of basis functions used in the decomposition.
@@
where $P(k)$ is the original primordial power spectrum. In practice we set $\phi_{\rm pk}=0$, and do not consider any phase offsets between the oscillations in the power spectrum and bispectrum. We use $k_{\star}=1\,\,h{\rm/Mpc}$ as with the bispectrum. Different physical processes can inject oscillations into the power spectrum, and various types have been studied with $N$-body simulations (e.g., Schaeffer & Schneider 2021; Stahl et al. 2025). This work is the first to simultaneously simulate resonance-driven features consistently in the power spectrum and bispectrum.
@@
Bispectra can also be generated if the ground state of the inflaton is excited, which is also referred to as a non-Bunch-Davies (NBD) vacuum (e.g., Chen et al. 2007a; Holman & Tolley 2008; Meerburg et al. 2009). The signal in these models generally peaks in the folded limit, $k_{1}+k_{2}\approx k_{3}$, though the exact behavior varies depeneding on the exact model under consideration. We only consider a subset of the NBD models studied in Planck Collaboration et al. (2014, 2016, 2020b) as the others were not accurately captured by our basis functions. Following the Planck nomenclature, we explore the “NBD mode 2” model, {mybox}
@@
Models with excited initial states are also often accompanied by resonant features. We use the template (Chen 2010b), {mybox}
@@
Finally, we consider the template from Dirac-Born-Infeld (DBI) inflation (Silverstein & Tong 2004; Alishahiha et al. 2004), {mybox}
@@ 3.1 Information in Weak Lensing Observations
In brief, we assume a survey footprint of 14,000 $\deg^{2}$ with a source galaxy number density of $n_{\rm gal}=30\,\,{\rm arcmin}^{-2}$, following the LSST Year 10 (Y10) prescription. We use five tomographic bins as in Anbajagane et al. (2023). Our summary statistics are the 2nd and 3rd moments of the lensing convergence field, $\kappa$. We measure all auto- and cross-correlations between the tomographic bins, and probe the scale dependence by smoothing the maps with 10 circular apertures between $3.2\hbox{${}^{\prime}$}<\theta<200\hbox{${}^{\prime}$}$. While our survey setup focuses on upcoming data (LSST Y10), these forecasts are still useful for existing datasets. A combination of lensing data from the Dark Energy Survey (Gatti et al. 2021; Secco et al. 2022; Amon et al. 2022) and the Dark Energy Camera All Data Everywhere (DECADE) project (Anbajagane et al. 2025a, d) gives one access to 270 million galaxies spanning 13,000 $\deg^{2}$ of the sky, which is close to the same area we consider here. These existing galaxy samples probe lower redshifts (Myles et al. 2021; Anbajagane et al. 2025b) than those anticipated from LSST Y10. However, our forecast here is still relevant as the signal is concentrated towards lower redshifts (Anbajagane et al. 2023, see their Figure 6).
@@
Figure 2 presents the forecasts on various bispectrum models for LSST Y10 lensing data. Each model is characterized by the amplitude $f_{\rm NL}$ and one or two additional parameters whose interpretations vary depending on the model. For each case, we generate simulations at selected values of the additional parameters, with the number of choices depending on the exact model, and compute the expected constraint on $f_{\rm NL}$ for those fixed parameter values. In Paper I, we quantified the competitiveness of our forecasts by comparing them to the Planck 2018 constraints of Sohn et al. (2024). While the models considered in this work similarly have constraints presented in Planck Collaboration et al. (2014, 2016, 2020b), these constraints are generally quoted in terms of signal-to-noise and not $\sigma(f_{\rm NL})$. As such, no direct comparison with our forecasted uncertainties is available. We therefore perform our own analysis by utilizing the templates of Figure 1 on the temperature and polarization data from Planck 2018.444Note that the combination of Planck 2018 with the upcoming Simons Observatory data is expected to improve constraints by 20-50% for the local, equilateral, and orthogonal templates (Ade et al. 2019). We can then expect a similar level of improvement, relative to the Planck 2018 constraints, for the templates studied here. This is done using the publicly available CMB-BEST pipeline (and associated data products). We detail the analysis choices further in Appendix B.
@@
As discussed in Paper I, our choice to perform single-parameter constraints—due to limitations in the simulation suite—means these results can be viewed as a best-case scenario when using the second and third moments. Anbajagane et al. (2023) show that marginalizing over additional parameters—such as $\sigma_{8}$, the root-mean squared amplitude of fluctuations at $z=0$ on scales of $8\,{\rm Mpc}/h$, and $\Omega_{m}$, the fraction of energy density contributed by matter, as well as two nuisance parameters related to the intrinsic alignment of galaxies (Troxel & Ishak 2015)—degrades constraints on $f_{\rm NL}$ by 20% to 30% for the standard $f_{\rm NL}$ templates (local, equilateral, and orthogonal). A similar degradation can be expected here as the templates we study are somewhat correlated with one or more of these standard templates. We note that existing works also limit their data constraints to a single-parameter fit when constraining $f_{\rm NL}$ (e.g., Planck Collaboration et al. 2020b; Sohn et al. 2024).
@@
Figure 3 presents the cross-correlation of the late-time signatures of different templates, as seen in the second and third moments of the lensing convergence field. We compute this by estimating the Fisher information of a two-model analysis, [^5] As discussed in Paper I, deviating from a single-parameter analysis can cause our estimates to be impacted by noise in the derivatives (Coulton & Wandelt 2023). We have confirmed that our estimates of the off-diagonal covariance changes by at most $\Delta{\rm corr}=0.3$ if we use one-third as many realizations. This does not change our qualitative discussions about the correlation between two templates. We do not make any quantitative statements. and evaluating the parameter covariance matrix. We normalize the off-diagonal component to obtain the correlation, and take the absolute value to isolate just the amplitude (and not sign) of the correlation. Values close to 1 indicate strong correlations between the late-time predictions of two templates, and those close to 0 indicate strong orthogonality between the same. The upper triangle shows results for datavectors in an LSST Y10 survey, while the lower triangle is the same for datavectors in a cosmic variance-limited survey, i.e., one with no shape noise in the lensing estimates. The black boxes approximately delineate different classes of the templates discussed in Section 2.2.

### v2

@@ 1 Introduction
Under the inflationary paradigm, the Universe underwent a rapid phase of expansion in its earlier moments (Guth 1981; Linde 1982; Guth 2004). This process also generates the initial fluctuations in the density field that subsequently grow and are observed in the cosmic microwave background and the large-scale structure. The fluctuations are commonly modeled to follow the statistics of a Gaussian random field. However, deviations from Gaussianity—called “Primordial Non-Gaussianities” (PNGs)—can arise from particle interactions during inflation (see Chen 2010b, for a review). Detecting such PNGs would provide a unique window into the microphysics of the early Universe, offering direct clues about the field content and interactions operative during inflation.
@@
There are also extensions to these standard scenarios that capture interactions of the inflaton with additional massless/massive particles. This framework, known as “cosmological collider physics” (Chen and Wang 2010; Baumann and Green 2012; Arkani-Hamed and Maldacena 2015; Lee et al. 2016; Arkani-Hamed et al. 2020), has recently been studied in the large-scale structure context by Goldstein et al. (2024) for the squeezed-limit bispectrum and by Anbajagane and Lee (2025, henceforth, Paper I) for the full bispectrum. In particular, in Paper I we introduced a method for generating initial density fields with arbitrary primordial bispectra. This enabled us to resolve, for the first time, over thirty collider PNG templates in the nonlinear regime. The inflationary signatures in structure formation from this regime are distinct from those signatures found on linear and quasi-linear scales, and therefore provide a complementary avenue for placing constraints on inflationary models (Anbajagane et al. 2023, Paper I). Furthermore, Paper I showed that the potential lensing-only constraints on such collider models can be competitive and complementary to the existing constraints from Planck 2018 (Sohn et al. 2024).
@@ 2.1 Generating Initial Conditions from Arbitrary Bispectra
To generate initial conditions with a given bispectrum, we follow the approach of Scoccimarro et al. (2012). This method requires the bispectrum (or kernel) to be separable, i.e., $B(k_{1},k_{2},k_{3})=f(k_{1})g(k_{2})h(k_{3})$. This condition is not always met for all bispectra considered in this work. Instead, we follow the approach of Paper I — which builds on the longstanding mode decomposition efforts in the CMB community (Komatsu et al. 2005; Smith and Zaldarriaga 2011; Fergusson et al. 2012; Sohn et al. 2023, 2024) and more recently in the large-scale structure community (e.g., Assassi et al. 2017; Lee and Dvorkin 2020; Chen et al. 2021) — by decomposing the bispectra into a sum of terms that are each individually factorizable as
@@ 2.2 Inflationary Models
We consider a wide variety of models following those explored in Planck Collaboration et al. (2014, 2016, 2020b), and choose a subset to simulate in our work. Many of these models, such as the scale-dependent ones or linear resonances, are manifestly separable (Münchmeyer et al. 2014). Others that are not manifestly separable due to a factor of $1/(k_{1}+k_{2}+k_{3})^{n}$, with $n>0$, can be approximated using the Schwinger parameterization (e.g., Smith and Zaldarriaga 2011). Thus, many bispectra also have template-specific approaches to decomposing them. Instead of following different methods tailored to each given bispectrum, we follow the Planck analyses in using mode decomposition methods on all templates, given we scan a wide range of model space. We now detail our chosen models below.
@@
We then consider templates corresponding to resonant particle interactions. In this case, the inflaton potential has a localized feature that generates oscillations in the bispectrum (see Achúcarro et al. 2022, for a review). The models used in Paper I also exhibit oscillations, but are derived from explicit couplings of the inflaton to other massive fields (e.g., Arkani-Hamed and Maldacena 2015), whereas the models we study here are described by phenomenological parameters alone. Sharp changes to the inflaton potential also induce oscillations at other orders, such as the power spectrum and trispectrum. In this work, we will first focus on oscillations only in the bispectrum, but then consider a class of models where the bispectrum and power spectrum both contain oscillations. We ignore the primordial trispectra as its impact on late-time structure is subdominant to the power spectra and bispectra. [^1] Note that there do exist some models where the primordial trispectrum can dominate over the bispectrum (Philcox 2025a, b). In general, the derivative of a signature in late-time structure, $X$, with the bispectrum amplitude $f_{\rm NL}$ is $\mathcal{O}(10^{3})$ higher than that with the trispectrum amplitude, $g_{\rm NL}$ (e.g., LoVerde and Smith 2011, see their ).
Oscillatory bispectra can be broadly classified into linear and logarithmic templates, given by (Chen et al. 2007b; Chen 2010b) {mybox}
@@
where $\omega$ is the frequency of the oscillations and $\phi=0$ is the phase. We explore a fairly small range in $\omega\in[0.1,2.5]$ relative to perturbation theory-based analyses of the CMB (e.g., Beutler et al. 2019b). Our choice is limited by the basis functions used in decomposing the templates. Oscillations of high order (large $w$) are not accurately reproduced by our existing basis functions. [^2] In principle, this can be alleviated by using an adequately high order for the Legendre polynomials in Eq. (2.3). However, doing so degrades the orthogonality of the basis set as discussed in Section 2.1 of Paper I. Higher values of $w$ could be explored by making improvements to the choice of basis functions used in the decomposition.
@@
where $P(k)$ is the original primordial power spectrum. In practice we set $\phi_{\rm pk}=0$, and do not consider any phase offsets between the oscillations in the power spectrum and bispectrum. We use $k_{\star}=1\,\,h{\rm/Mpc}$ as with the bispectrum. Different physical processes can inject oscillations into the power spectrum, and various types have been studied with $N$-body simulations (e.g., Schaeffer and Schneider 2021; Stahl et al. 2025). This work is the first to simultaneously simulate resonance-driven features consistently in the power spectrum and bispectrum.
@@
Bispectra can also be generated if the ground state of the inflaton is excited, which is also referred to as a non-Bunch-Davies (NBD) vacuum (e.g., Chen et al. 2007a; Holman and Tolley 2008; Meerburg et al. 2009). The signal in these models generally peaks in the folded limit, $k_{1}+k_{2}\approx k_{3}$, though the exact behavior varies depeneding on the exact model under consideration. We only consider a subset of the NBD models studied in Planck Collaboration et al. (2014, 2016, 2020b) as the others were not accurately captured by our basis functions. Following the Planck nomenclature, we explore the “NBD mode 2” model, {mybox}
@@
Models with excited initial states are also often accompanied by resonant features. We use the template (Chen 2010a), {mybox}
@@
Finally, we consider the template from Dirac-Born-Infeld (DBI) inflation (Silverstein and Tong 2004; Alishahiha et al. 2004), {mybox}
@@ 3.1 Information in Weak Lensing Observations
Measurements of weak lensing on smaller scales have been shown to have a strong sensitivity to $f_{\rm NL}$ (Shirasaki et al. 2012; Anbajagane et al. 2023; Anbajagane and Lee 2025). This is primarily because $f_{\rm NL}$ alters the formation and evolution of massive structures, which dominate the signals on such scales. Conceptually, this signal can be viewed as originating from the impact of $f_{\rm NL}$ on the halo mass function (HMF), which has been extensively documented by the community (Dalal et al. 2008; Shirasaki et al. 2012; Marian et al. 2011; Hilbert et al. 2012; Jung et al. 2023). However, a more precise statement is that the signal originates from changes to the statistics of peaks in the density field, which have a clear dependence of $f_{\rm NL}$ as detailed in Dalal et al. (2008, see their Appendix). See Section 5 of Anbajagane et al. (2023) for more detail on the origin of the $f_{\rm NL}$ signal in weak lensing.
@@
In brief, we assume a survey footprint of 14,000 $\deg^{2}$ with a source galaxy number density of $n_{\rm gal}=30\,\,{\rm arcmin}^{-2}$, following the LSST Year 10 (Y10) prescription. We use five tomographic bins as in Anbajagane et al. (2023). Our summary statistics are the 2nd and 3rd moments of the lensing convergence field, $\kappa$. We measure all auto- and cross-correlations between the tomographic bins, and probe the scale dependence by smoothing the maps with 10 circular apertures between $3.2\hbox{${}^{\prime}$}<\theta<200\hbox{${}^{\prime}$}$. While our survey setup focuses on upcoming data (LSST Y10), these forecasts are still useful for existing datasets. A combination of lensing data from the Dark Energy Survey (Gatti et al. 2021; Secco et al. 2022; Amon et al. 2022) and the Dark Energy Camera All Data Everywhere (DECADE) project (Anbajagane et al. 2025e, c, d) gives one access to 270 million galaxies spanning 13,000 $\deg^{2}$ of the sky, which is close to the same area we consider here. These existing galaxy samples probe lower redshifts (Myles et al. 2021; Anbajagane et al. 2025a) than those anticipated from LSST Y10. However, our forecast here is still relevant as the signal is concentrated towards lower redshifts (Anbajagane et al. 2023, see their Figure 6).
The Fisher estimation requires two inputs: the covariance of our measured datavector and the derivative of the datavector with respect to the parameter of interest ($f_{\rm NL}$). We estimate the former with 6000 independent realizations of the simulated datavector, leading to a Kaufman-Hartlap factor of $\gtrsim 0.95$ (Kaufman 1967; Hartlap et al. 2007) for our analysis, while the latter is estimated using 30 independent realizations. Derivatives contaminated by numerical noise can artificially break parameter degeneracies and thereby overestimate constraining power (Coulton and Wandelt 2023). In our case, we limit our key analyses to a single-parameter forecast on just $f_{\rm NL}$. [^4] Other approaches compress the datavector into a low-dimensional space where the impact of noise on the Fisher information is far less pronounced (Coulton and Wandelt 2023). However, defining the compression requires simulations as well, and given we only have 30 realizations for the $f_{\rm NL}$ derivatives, we do not pursue this approach. As a result, there is no such degeneracy breaking and the impact of numerical noise is subdued. We have explicitly confirmed our results are converged at the 2-3% level if we estimate derivatives using half the available realizations.
@@
Figure 2 presents the forecasts on various bispectrum models for LSST Y10 lensing data. Each model is characterized by the amplitude $f_{\rm NL}$ and one or two additional parameters whose interpretations vary depending on the model. For each case, we generate simulations at selected values of the additional parameters, with the number of choices depending on the exact model, and compute the expected constraint on $f_{\rm NL}$ for those fixed parameter values. In Paper I, we quantified the competitiveness of our forecasts by comparing them to the Planck 2018 constraints of Sohn et al. (2024). While the models considered in this work similarly have constraints presented in Planck Collaboration et al. (2014, 2016, 2020b), these constraints are generally quoted in terms of signal-to-noise and not $\sigma(f_{\rm NL})$. As such, no direct comparison with our forecasted uncertainties is available. We therefore perform our own analysis by utilizing the templates of Figure 1 on the temperature and polarization data from Planck 2018.555Note that the combination of Planck 2018 with the upcoming Simons Observatory data is expected to improve constraints by 20-50% for the local, equilateral, and orthogonal templates (Ade et al. 2019). We can then expect a similar level of improvement, relative to the Planck 2018 constraints, for the templates studied here. This is done using the publicly available CMB-BEST pipeline (and associated data products). We detail the analysis choices further in Appendix B.
@@
We note that our above statement, that structure-based constraints could outperform the CMB for signals peaking on small scales, is specific to the lensing analysis being discussed here. Conventional structure-based analyses of $f_{\rm NL}$ (e.g., Cabass et al. 2022a; Philcox et al. 2022; D’Amico et al. 2022) focus on the galaxy density field and use perturbative approaches to model both the galaxy bias and the spatial correlations in the matter density field. In this case there is a large uncertainty in (analytically) modeling the non-linear gravitational evolution of structure, and this significantly reduces the potential constraining power from using more non-linear scales. Such perturbative analyses are often limited to the same scales probed by CMB datasets, and therefore will not offer the same precision improvements discussed above. In the case of lensing, we only need to model the matter density field and therefore use simulation-based models, which can probe these smaller scales without incurring a significant penalty on constraining power. In principle, analyses of the galaxy density field can also avoid this penalty using simulations. However, the simulations needed for galaxy analyses are far more computationally expensive compared to their lensing counterparts (Anbajagane et al. 2023, see their Section 5.1).
@@
As discussed in Paper I, our choice to perform single-parameter constraints—due to limitations in the simulation suite—means these results can be viewed as a best-case scenario when using the second and third moments. Anbajagane et al. (2023) show that marginalizing over additional parameters—such as $\sigma_{8}$, the root-mean squared amplitude of fluctuations at $z=0$ on scales of $8\,{\rm Mpc}/h$, and $\Omega_{m}$, the fraction of energy density contributed by matter, as well as two nuisance parameters related to the intrinsic alignment of galaxies (Troxel and Ishak 2015)—degrades constraints on $f_{\rm NL}$ by 20% to 30% for the standard $f_{\rm NL}$ templates (local, equilateral, and orthogonal). A similar degradation can be expected here as the templates we study are somewhat correlated with one or more of these standard templates. We note that existing works also limit their data constraints to a single-parameter fit when constraining $f_{\rm NL}$ (e.g., Planck Collaboration et al. 2020b; Sohn et al. 2024).
@@
Figure 3 presents the cross-correlation of the late-time signatures of different templates, as seen in the second and third moments of the lensing convergence field. We compute this by estimating the Fisher information of a two-model analysis, [^6] As discussed in Paper I, deviating from a single-parameter analysis can cause our estimates to be impacted by noise in the derivatives (Coulton and Wandelt 2023). We have confirmed that our estimates of the off-diagonal covariance changes by at most $\Delta{\rm corr}=0.3$ if we use one-third as many realizations. This does not change our qualitative discussions about the correlation between two templates. We do not make any quantitative statements. and evaluating the parameter covariance matrix. We normalize the off-diagonal component to obtain the correlation, and take the absolute value to isolate just the amplitude (and not sign) of the correlation. Values close to 1 indicate strong correlations between the late-time predictions of two templates, and those close to 0 indicate strong orthogonality between the same. The upper triangle shows results for datavectors in an LSST Y10 survey, while the lower triangle is the same for datavectors in a cosmic variance-limited survey, i.e., one with no shape noise in the lensing estimates. The black boxes approximately delineate different classes of the templates discussed in Section 2.2.